# End-to-End and Highly-Efficient Differentiable Simulation for Robotics

### v1

@@ 3.3 Efficient computation: exploiting kinematic-induced sparsity
where $B$ is block diagonal with identity for blocks of $\mathcal{A}_{\text{stk}}$ and the basis change $R^{\top}P$ for blocks of $\mathcal{A}_{\text{sld}}$; the complete construction is given in the Appendix.
Computing the right-hand side of (10) requires evaluating $\frac{\mathrm{d}G}{\text{d}\theta}\bm{\lambda}^{*}+\frac{\mathrm{d}g}{\text{d}\theta}$. By recalling that $G\bm{\lambda}^{*}+g=J_{c}\bm{v}^{+}$ is the contact point velocity, this term exactly corresponds to the derivatives wrt $\theta$ of the contact point velocity with $\bm{\lambda}^{*}$ taken constant. We have
@@
The first term is already computed thanks to the ABA derivatives (6[34], and the second term $\left.\frac{\text{d}J_{c}\bm{v}^{+}}{\text{d}\theta}\right|_{v=v^{+}}$, which is the derivatives of the contact velocity with $\bm{v}^{+}$ assumed to be constant, can be computed at a reduced cost via the partial derivatives of the forward kinematics [34] evaluated in $\bm{q},\bm{v}^{+}$. This allows us to avoid the expensive computation of $\frac{\mathrm{d}G}{\text{d}\theta}$ as it is a tensor in general, while only its product with $\bm{\lambda}$ is required. It is worth noting at this stage that the term $\left.\frac{\text{d}J_{c}\bm{v}^{+}}{\text{d}\theta}\right|_{v=v^{+}}$ also depends on the geometry of the contact. Therefore, computing gradients w.r.t. to the configuration $\bm{q}$ requires evaluating an additional term for the variations of $J_{c}$ induced by the variations of contact points on the local geometries [22] and presented next.
@@
3.4 Collision detection contribution
The collision detection phase depends on the body poses induced by the configuration $\bm{q}$. Thus, when $\theta$ depends on $\bm{q}$, one must consider the variation of the contact given variations of $\bm{q}$. Recent work on differentiable collision detection [22, 21] allows computing the derivative of the normal and contact point w.r.t. the poses of the bodies.
By choosing a function that constructs a contact frame $c$ from a contact point and its normal, we compute the derivative of this frame w.r.t. the body poses. Chaining this derivative with the usual kinematics Jacobian, which relates the variation of $\bm{q}$ to the variation of body poses, one can obtain $\frac{\text{d}c}{\text{d}\theta}$.
The frame $c$ intervenes in $J_{c}$ through a change of frame (the adjoint of the placement). By leveraging spatial algebra[30] (see the appendix for the details), we calculate $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}$ and $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}$. We add the first collision term $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ in $\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}$ and the second collision term $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ in $\left.\frac{\text{d}J_{c}\bm{v}^{+}}{\text{d}\theta}\right|_{v=v^{+}}$ to account for the variation of $J_{c}$ due to the variation of the contact points. The complete details of the terms and simulation derivative $\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}$ are reported in the appendix.
4 Experiments
In this section, we first demonstrate state-of-the-art computational timings when computing simulation gradients for various robotic systems (Fig. 2). Second, we apply our approach to solve two inverse problems involving contact dynamics: retrieving an initial condition that leads to a target final state and finding a torque that yields a null acceleration on a quadruped. Additional experiments are available in the Appendix.
@@ 4 Experiments
Implementation details. We have implemented our analytical derivatives in C++ for efficiency. We leverage open-source software of the community: Eigen [39] for efficient linear algebra, Pinocchio [35] for fast rigid body dynamics and their derivatives, and HPP-FCL [40] for high-speed collision detection. The code associated with this paper will be released as open-source as part of the Simple simulator. All the experiments are performed on a single core of an Apple M3 CPU.
4.1 Timings
Half-cheetah Humanoid UR5 Go1 H1 Atlas Number of DoFs 12 27 7 18 25 36 Number of geometries 9 20 9 39 25 89 Number of collision pairs 27 161 26 494 255 3399 Number of vertices per mesh N/A N/A 200 N/A 700 N/A
Table 2: Details of scenarios for computational timings. For the number of vertices per mesh, N/A indicates a scenario which does not contain any mesh.
@@ 4.1 Timings
Table 3: Comparative analysis between ours, MuJoCo, and MJX frameworks. Timing statistics (mean and standard deviation in microseconds) for simulation, gradient, and finite-differences computation for one simulation step. For MJX, N/A denotes scenarios where geometries were not supported.
The computational efficiency of our approach is evaluated by measuring the average time required to compute the full Jacobian of the simulator, i.e., $\frac{\text{d}\bm{v}^{+}}{\text{d}\bm{q},\bm{v},\bm{\tau}}$, along a trajectory. We consider various scenarios ranging from simple systems composed of basic geometry primitives (MuJoCo’s half-cheetah and humanoid) to more complex and realistic robots with multiple DoFs and complex geometries (UR5, Unitree Go1, and Boston Dynamics Atlas). Tab. 2 reports the numbers associated with the different robots considered. To stabilize the simulation behaviors, we compute contact collision patches (composed of 4 contact points each), thus substantially increasing the dimensions of the problem to solve.
Tab. 3 demonstrates computational timings for gradient computation that are of the same order of magnitude as simulation and significantly faster than central finite differences. As another point of comparison, Nimble [10] requires 1ms and 16ms on half-cheetah and Atlas, corresponding to an approximate speedup factor of 100 for our method. In Tab. 3, we also compare our approach to MuJoCo MJX GPU simulation pipeline and find it to be competitive even though it operates on a single CPU core. Importantly, our performance gain is obtained although we work on the unrelaxed NCP and with full meshes descriptions (cf. Tab. 1).
4.2 Inverse problems
@@
where $\bm{q}_{T}$ is the final configuration and depends on the initial velocity and impulse. Our forward-mode differentiation allows us to efficiently compute the Jacobian of $\bm{q}_{T}$ wrt $\bm{\theta}$. We leverage this feature to implement a Gauss-Newton (GN) approximation of the Hessian of (12). Fig. 3 demonstrates the benefits of using our implicit gradients over finite-differences in order to reach a precise solution of (12). Moreover, exploiting the full Jacobian in a quasi-Newton algorithm also reduces the number of iterations compared to a vanilla gradient descent.
@@
$\min_{\bm{\tau}_{\text{act}}}\frac{1}{2}\left\|\bm{v}^{+}(\bm{q},\bm{v},S^{T}\bm{\tau}_{\text{act}})-\bm{v}^{*}\right\|_{2}^{2},$ (13)
@@
Figure 4: Contact inverse dynamics on an underactuated Go1 quadruped can be efficiently performed via a Gauss-Newton algorithm by using Jacobians of our differentiable physics engine.
5 Limitations
@@
This paper introduces an end-to-end differentiable physics pipeline for robotics based on the implicit differentiation of the non-relaxed NCP for contacts. By avoiding any relaxation, we prevent the appearance of unphysical simulation artifacts. Moreover, exploiting the sparsity induced by the robot kinematic chains and leveraging the derivatives of rigid body algorithms allows us to achieve state-of-the-art timings, with a speed-up of at least $100$ compared to alternative solutions of the state of the art. In an MPC context where the dynamics and its derivatives are evaluated at a high frequency, the gains in physical realism and efficiency could determine the controller’s overall performance.
@@
Yet, as the NCP induces inherently non-smooth dynamics, exploiting its gradients requires dedicated algorithms when addressing downstream optimization tasks. Some previous works [41, 42, 37] leverage randomized smoothing techniques that provide smooth gradient estimates from simulation and gradient samples. Alternative solutions relax the physics, either explicitly [2, 1, 43] or implicitly, by leveraging interior-points (IP) methods [13].
@@
Conclusion. In future work, we plan to integrate our qualitative gradient approach with recent advancements in control frameworks to tackle more complex robotics tasks. Specifically, this paper introduces an efficient method for computing physics gradients, which paves the way for first-order policy learning algorithms such as SHAC [45]. It is also worth noticing that our approach is not limited to rigid-body robots, but could also be leveraged for soft dynamics in general to design and control soft robots [46] and could be adapted to use implicit integrator as in [47]. The proposed differentiable simulation of an unrelaxed physics model is a crucial step toward reducing the Sim2Real gap [48], and future work may adapt learning algorithms to effectively achieve this goal. Finally, we hope this work will serve as a catalyzer in the robotics and learning communities and motivate the development of new reinforcement learning and trajectory optimization methods leveraging simulation gradients in order to accelerate the discovery of complex robot movements in contact.
Acknowledgments
This work was supported in part by the French government under the management of Agence Nationale de la Recherche (ANR) as part of the ”Investissements d’avenir” program, references ANR-19-P3IA-0001 (PRAIRIE 3IA Institute) and ANR-22-CE33-0007 (INEXACT), the European project AGIMUS (Grant 101070165), the Louis Vuitton ENS Chair on Artificial Intelligence and the Casino ENS Chair on Algorithmic and Machine Learning. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funding agencies.
Appendix
This appendix document reports the details associated with the submission entitled End-to-End and Highly-Efficient Differentiable Simulation for Robotics. In particular, we provide detailed equations related to the three modes of the NCP implicit differentiation, the contributions of collision detection, and the detailed expression of multibody derivatives accounting for contact and collision derivatives.
Appendix A - Sliding mode
In this subsection, we detail the equations (Eq. (9) of the paper) of the sensitivity analysis of the contact force in the case of a sliding mode ($\|\bm{\sigma}_{T}\|>0$).
In sliding mode, both the contact forces and the contact point velocity are on the border of their cone. Thus, from the NCP we have:
@@ Acknowledgments
$A\mathrm{d}\bm{\lambda}=-\frac{1}{\alpha}Hd\bm{\sigma}_{T}$ (16a)
@@
Because $\bm{\lambda}$ is constrained to stay on the boundary of the cone, its variations $\text{d}\bm{\lambda}$ live in the tangent 2D plane whose basis is $R=\begin{pmatrix}\frac{\bm{\lambda}}{\|\bm{\lambda}\|}&\bm{e}_{z}\times\frac{\bm{\sigma}_{T}}{\|\bm{\sigma}_{T}\|}\end{pmatrix}\in\mathbb{R}^{3\times 2}$. Applying the change of variable $\mathrm{d}\bm{\lambda}=R\mathrm{d}\tilde{\bm{\lambda}}$ and multiplying the previous system 18 by $R^{\top}$ allows getting a linear system of reduced dimension:
@@
Appendix B - Implicit NCP gradient system
In this subsection, we detail the final system solved to compute the gradients of the NCP (equation (10) of the paper).
Following the paper notations, we denote $\mathcal{A}_{\text{brk}}$, $\mathcal{A}_{\text{stk}}$, and $\mathcal{A}_{\text{sld}}$ as the sets of contact indices corresponding to the breaking, sticking, and sliding contacts respectively and $n_{\text{brk}}$, $n_{\text{stk}}$, and $n_{\text{sld}}$ their cardinals and $n=n_{\text{brk}}+n_{\text{stk}}+n_{\text{sld}}$ the total number of contacts. For the implicit gradients system, we removed the $\text{d}\bm{\lambda}$ variables associated with contacts in $\mathcal{A}_{brk}$ and sorted the remaining ones by putting first the components associated with the contacts in $\mathcal{A}_{stk}$ before the reduced ones $\text{d}\tilde{\bm{\lambda}}$ of $\mathcal{A}_{sld}$. Then the total variation of $\bm{\lambda}$ is $\text{d}\bm{\lambda}=CX\in\mathbb{R}^{3n}$ with

### v2

@@ III-C Efficient computation: exploiting kinematic-induced sparsity
where $B$ is block diagonal with identity for blocks of $\mathcal{A}_{\text{stk}}$ and the basis change $R^{\top}P$ for blocks of $\mathcal{A}_{\text{sld}}$; the complete construction is given in the Appendix B.
Computing the right-hand side of (11) requires evaluating the derivative of $G\bm{\lambda}^{*}+g$ with $\bm{\lambda}^{*}$ taken constant:
$\frac{\mathrm{d}G}{\text{d}\theta}\bm{\lambda}^{*}+\frac{\mathrm{d}g}{\text{d}\theta}=\left.\frac{\text{d}G\bm{\lambda}^{*}+g}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}.$ (12)
By recalling that $G\bm{\lambda}^{*}+g=J_{c}\bm{v}^{+}$ (see Eq. 56)is the contact point velocity, this term exactly corresponds to the derivatives w.r.t. $\theta$ of the contact point velocity with $\bm{\lambda}^{*}$ taken constant. Calculating the derivative, we obtain:
@@
The first term is already computed thanks to the ABA derivatives (7[6], and the second term $\left.\frac{\text{d}J_{c}\bm{v}^{+}}{\text{d}\theta}\right|_{v=v^{+}}$, which is the derivatives of the contact velocity with $\bm{v}^{+}$ assumed to be constant, can be computed at a reduced cost via the partial derivatives of the forward kinematics [6] evaluated in $\bm{q},\bm{v}^{+}$. This allows us to avoid the expensive computation of $\frac{\mathrm{d}G}{\text{d}\theta}$ as it is a tensor in general, while only its product with $\bm{\lambda}$ is required. It is worth noting at this stage that the term $\left.\frac{\text{d}J_{c}\bm{v}^{+}}{\text{d}\theta}\right|_{v=v^{+}}$ also depends on the geometry of the contact. Therefore, computing gradients w.r.t. to the configuration $\bm{q}$ requires evaluating an additional term for the variations of $J_{c}$ induced by the variations of contact points on the local geometries [33] and presented in the next subsection.
@@
III-D Collision detection contribution
The collision detection phase depends on the body poses induced by the configuration $\bm{q}$. Thus, when $\theta$ depends on $\bm{q}$, one must consider the variation of the contact location given variations of $\bm{q}$. Recent work on differentiable collision detection [33, 43] allows computing the derivative of the normal and contact point w.r.t. the poses of the bodies.
By choosing a function that constructs a contact frame $c$ from a contact point and its normal, we compute the derivative of this frame w.r.t. the body poses. Chaining this derivative with the usual kinematics Jacobian, which relates the variation of $\bm{q}$ to the variation of body poses, one can obtain $\frac{\text{d}c}{\text{d}\theta}$. The frame $c$ intervenes in $J_{c}$ through a change of frame (the adjoint of the placement). By leveraging spatial algebra[16] (see Appendix C for the details), we calculate $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}$ and $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}$. We add the first collision term $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ in $\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}$ and the second collision term $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ in $\left.\frac{\text{d}J_{c}\bm{v}^{+}}{\text{d}\theta}\right|_{v=v^{+}}$ to account for the variation of $J_{c}$ due to the variation of the contact points.
The complete details of the terms and simulation derivative $\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}$ are reported in Appendix D.
IV Experiments
In this section, we first demonstrate state-of-the-art computational timings when computing simulation gradients for various robotic systems (Fig. 2). Second, we apply our approach to solve two inverse problems involving contact dynamics: retrieving an initial condition that leads to a target final state and finding a torque that yields a null acceleration on a quadruped. Eventually, we use our approach in first-order policy learning algorithms in order to efficiently train control policies on various systems. The experiments as well as additional experiments presented in Appendix E are available in the video attached.
@@ IV Experiments
Implementation details. We have implemented our analytical derivatives in C++ for efficiency. We leverage open-source software of the community: Eigen [22] for efficient linear algebra, Pinocchio [7] for fast rigid body dynamics and their derivatives, and HPP-FCL [35] for high-speed collision detection. The code associated with this paper will be released as open-source upon acceptance. All the experiments are performed on a single core of an Apple M3 CPU.
IV-A Timings
@@
TABLE II: Details of scenarios for computational timings. For the number of vertices per mesh, N/A indicates a scenario that does not contain any mesh.
[…]
TABLE III: Comparative analysis between ours, MuJoCo, and MJX frameworks. Timing statistics (mean and standard deviation in microseconds) for simulation, gradient, and finite-differences computation for one simulation step. For MJX, N/A denotes scenarios where geometries were not supported.
The computational efficiency of our approach is evaluated by measuring the average time required to compute the full Jacobian of the simulator, i.e., $\frac{\text{d}\bm{v}^{+}}{\text{d}\bm{q},\bm{v},\bm{\tau}}$, along a trajectory. We consider various scenarios ranging from simple systems composed of basic geometry primitives (MuJoCo’s half-cheetah and humanoid) to more complex and realistic robots with multiple DoFs and complex geometries (UR5, Unitree Go1, and Boston Dynamics Atlas). Tab. II reports the numbers associated with the different robots considered. To stabilize the simulation behaviors, we compute contact collision patches (composed of 4 contact points each), thus substantially increasing the dimensions of the problem to solve.
Tab. III demonstrates computational timings for gradient computation that are of the same order of magnitude as simulation and significantly faster than central finite differences. As another point of comparison, Nimble [46] requires 1ms and 16ms on half-cheetah and Atlas, corresponding to an approximate speedup factor of 100 for our method. In Tab. III, we also compare our approach to MuJoCo MJX GPU simulation pipeline. The numbers for gradients computed on GPU correspond to the samples generated after one second when using all the threads of a Nvidia A100. We find our approach to be competitive even though it operates on a single CPU core. Importantly, our performance gain is obtained although we work on the unrelaxed NCP and with full meshes descriptions (cf. Tab. I). In general, this is not achieved by current GPU simulation approaches which have to operate on relaxed physical principles and simplified geometries due to hardware constraints.
IV-B Inverse problems
@@
where $\bm{q}_{T}$ is the final configuration and depends on the initial velocity and impulse. Our forward-mode differentiation allows us to efficiently compute the Jacobian of $\bm{q}_{T}$ w.r.t. $\bm{\theta}$. We leverage this feature to implement a Gauss-Newton (GN) approximation of the Hessian of (14). Fig. 3 demonstrates the benefits of using our implicit gradients over finite-differences in order to reach a precise solution of (14). Moreover, exploiting the full Jacobian in a quasi-Newton algorithm also reduces the number of iterations compared to a vanilla gradient descent.
@@
$\min_{\bm{\tau}_{\text{act}}}\frac{1}{2}\left\|\bm{v}^{+}(\bm{q},\bm{v},S^{\top}\bm{\tau}_{\text{act}})-\bm{v}^{*}\right\|_{2}^{2},$ (15)
@@
Figure 4: Contact inverse dynamics on an underactuated Go1 quadruped can be efficiently performed via a Gauss-Newton algorithm by leveraging the derivatives of our differentiable simulator.
IV-C Applications to Policy Learning
Model-free reinforcement learning relies on zeroth-order gradient estimation via the policy gradient theorem, often leading to high variance and slow convergence. In contrast, first-order gradient-based optimization using analytical gradients from a differentiable simulator enables more efficient policy updates e.g. SHAC [48] and AHAC [19].
Figure 5: SHAC vs. PPO on cartpole swing up (upper) and hopper (lower). SHAC algorithm leverages differentiable simulation to achieve improved sample efficiency.
We integrate our differentiable simulator with the first-order, on-policy algorithm SHAC and compare it to the zeroth-order, on-policy algorithm PPO. We evaluate two illustrative systems: the cartpole swing-up and the MuJoCo hopper. The performances are reported in Fig.  5. While these systems are relatively small, they are not trivial due to their non-smooth dynamics. The cartpole involves joint limits, and the hopper combines joint constraints with ground contact interactions, making these tasks representative of the complexities encountered in real-world scenarios. Commonly used robotics benchmarks in RL are from Gymnasium and rely on the MuJoCo simulator [42], which relaxes contact constraints using compliance [30]. First-order RL benchmarks with SHAC or AHAC work with even more relaxed dynamics by using the dflex simulator. The later models all constraints with a spring-damper approximation which results in smoother dynamics. In contrast, our simulation approach allows modeling non-compliant contacts and joint limits without any such relaxations [8]. In turn, the absence of relaxation directly affects the smoothness and continuity of the gradients, often making it challenging to leverage them for gradient based optimization. Certain reward terms in Gymnasium environments are non-smooth and require smoothing, as proposed in [48], to allow informative gradient computation (see section F in the Appendix).
The experiments show that using our differentiable simulator with SHAC leads to greater sample efficiency compared to PPO. However, the training process is noticeably noisier. This may arise from the stiffness of our simulator, which models non-smooth constraints directly rather than using the commonly employed relaxations or spring-damper systems. Although differentiable simulation reduces gradient variance, the resulting gradients can become highly unstable in stiff scenarios. Mitigating this instability is a challenging problem [19] and may require effective gradient smoothing strategies. Potential approaches include adaptive scheduling of a compliance term for the constraints or leveraging solver methods that inherently apply smoothing, such as interior point solvers [37, 24], to enhance stability without sacrificing efficiency.
V Limitations
@@
This paper introduces an end-to-end differentiable physics pipeline for robotics based on the implicit differentiation of the non-relaxed NCP for contacts. By avoiding any relaxation, we prevent the appearance of nonphysical simulation artifacts. Moreover, exploiting the sparsity induced by the robot kinematic chains and leveraging the derivatives of rigid body algorithms allows us to achieve state-of-the-art timings, with a speed-up of at least $100$ compared to alternative solutions of the state of the art. In an MPC context where the dynamics and its derivatives are evaluated at a high frequency, the gains in physical realism and efficiency could determine the controller’s overall performance.
@@
Yet, as the NCP induces inherently non-smooth dynamics, exploiting its gradients requires dedicated algorithms when addressing downstream optimization tasks. Our experiments with first-order reinforcement learning highlight this challenge: directly using the simulator’s raw analytical gradients with NCP-modeled constraints results in efficient but sometimes less stable convergence. Some previous works [40, 31, 36] leverage randomized smoothing techniques that provide smooth gradient estimates from simulation and gradient samples. Alternative solutions relax the physics, either explicitly [42, 34, 28] or implicitly, by leveraging interior-points (IP) methods [24].
Until now, none of the previous first-order RL works [48, 19] showed a sim-to-real transfer on real robotic hardware. This might be due to the significant difference between the smoothed simulators and reality. Therefore, working with the challenging-to-use but more accurate non-smooth gradients from more accurate simulators might pave the way to a sim-to-real transfer.
@@
VI Conclusion
In future work, we plan to combine our qualitative gradient approach with smoothing techniques to ease the integration within recent control frameworks to tackle more complex robotics tasks. Specifically, this paper introduces an efficient method for computing physics gradients, which are used in first-order policy learning algorithms such as SHAC [48] or AHAC[19]. The next step is extending our SHAC experiments to more complex real-world robotics systems. It is also worth noticing that our approach is not limited to rigid-body robots, but could also be leveraged for soft dynamics in general to design and control soft robots [13] and could be adapted to use implicit integrator as in [9]. The proposed differentiable simulation of an unrelaxed physics model is a crucial step toward reducing the Sim2Real gap [26], and future work may adapt learning algorithms to effectively achieve this goal. Finally, we hope this work will serve as a catalyzer in the robotics and learning communities and motivate the development of new reinforcement learning and trajectory optimization methods leveraging simulation gradients in order to accelerate the discovery of complex robot movements in contact.
Appendix A Sliding mode
In this section, we detail the equations (Eq. 10 of the paper) of the sensitivity analysis of the contact force in the case of a sliding mode ($\|\bm{\sigma}_{T}\|>0$). In sliding mode, both the contact forces and the contact point velocity are on the border of their cone. Thus, from the NCP we have:
@@ Appendix A Sliding mode
$K\mathrm{d}\bm{\lambda}=-\frac{1}{\alpha}Hd\bm{\sigma}_{T}$ (18a)
@@
Because $\bm{\lambda}$ is constrained to stay on the boundary of the cone, its variations $\text{d}\bm{\lambda}$ live in the tangent 2D plane whose basis is $R=\begin{pmatrix}\frac{\bm{\lambda}}{\|\bm{\lambda}\|}&\bm{e}_{z}\times\frac{\bm{\sigma}_{T}}{\|\bm{\sigma}_{T}\|}\end{pmatrix}\in\mathbb{R}^{3\times 2}$. Applying the change of variable $\mathrm{d}\bm{\lambda}=R\mathrm{d}\tilde{\bm{\lambda}}$ and multiplying the previous system 20 by $R^{\top}$ allows getting a linear system of reduced dimension:
@@
Appendix B Implicit NCP gradient system
In this section, we detail the final system solved to compute the gradients of the NCP (equation (10) of the paper). Following the paper notations, we denote $\mathcal{A}_{\text{brk}}$, $\mathcal{A}_{\text{stk}}$, and $\mathcal{A}_{\text{sld}}$ as the sets of contact indices corresponding to the breaking, sticking, and sliding contacts respectively and $n_{\text{brk}}$, $n_{\text{stk}}$, and $n_{\text{sld}}$ their cardinals and $n=n_{\text{brk}}+n_{\text{stk}}+n_{\text{sld}}$ the total number of contacts. For the implicit gradients system, we removed the $\text{d}\bm{\lambda}$ variables associated with contacts in $\mathcal{A}_{brk}$ and sorted the remaining ones by putting first the components associated with the contacts in $\mathcal{A}_{stk}$ before the reduced ones $\text{d}\tilde{\bm{\lambda}}$ of $\mathcal{A}_{sld}$. Then the total variation of $\bm{\lambda}$ is $\text{d}\bm{\lambda}=CX\in\mathbb{R}^{3n}$ with